OpenStudy (anonymous):

URGENT: PLEASE HELP Solve the equation. (Enter your answers as a comma-separated list.) 36^(x − 7) = 216^(4 − x) x =

OpenStudy (anonymous):

how urgent?

OpenStudy (anonymous):

haha yay youre back

OpenStudy (anonymous):

30 minutes

OpenStudy (anonymous):

ugent, hw is due at 11

OpenStudy (anonymous):

\[36=6^2\\ 216=6^3\]

OpenStudy (anonymous):

that means you can write these with the same base \[6^{2x-14}=6^{12-3x}\]

OpenStudy (anonymous):

all that is left for you to do is solve \[2x-14=12-3x\]

OpenStudy (anonymous):

let me know when you get \[\frac{26}{5}\]

OpenStudy (anonymous):

whats going on :(

OpenStudy (anonymous):

it wouldnt let me post

OpenStudy (anonymous):

i got those two

OpenStudy (anonymous):

did you see my answer?

OpenStudy (anonymous):

i had like 2x=26-3x

OpenStudy (anonymous):

26/5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

24 minutes to spare !

OpenStudy (anonymous):

thats the final answer??

OpenStudy (anonymous):

yes unless it wants a decimal

OpenStudy (anonymous):

yes 24 minutes but i have another 2 questions im stuck on :(

OpenStudy (anonymous):

post lets knock 'em the hell out

OpenStudy (anonymous):

Okay thank you! theyre not long

OpenStudy (anonymous):

that was right by the way

OpenStudy (anonymous):

A pendulum in a grandfather clock is 3 feet long and swings back and forth along a 4-inch arc. Approximate the angle (in degrees) through which the pendulum passes during one swing. (Round your answer to three decimal places.)

OpenStudy (anonymous):



OpenStudy (anonymous):